OpenStudy (anonymous):

Secx-(1/secx)=sinx tanx

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf sec(x)-\frac{1}{sec(x)} =sin(x)tan(x) }\) \(\Large\color{red}{ \bf \frac{sec^2(x)}{sec(x)}-\frac{1}{sec(x)} =sin(x)tan(x) }\) \(\Large\color{red}{ \bf \frac{tan^2(x)}{sec(x)} =sin(x)tan(x) }\) \(\Large\color{red}{ \bf \frac{tan(x)}{sec(x)} =sin(x) }\) \(\Large\color{red}{ \bf tan(x)cos(x)=sin(x) }\)

OpenStudy (neer2890):

\[\frac{ 1 }{ \cos x}- \cos x=\frac{ 1-\cos ^{2}x }{\cos x }= \frac{ \sin ^{2}x }{ \cos x }\] sinx sinx/cosx= sinx tan x